solve x^2+4x+4=0
pls andwer fast
update your webpage
 
 

Grade:11

1 Answers

Pallav Panwar
6 years ago
x^2+4x+4=0So determinant of the equation= b^2-4ac=0Hence the equation will have two equal roots.x= (-b +(b^2-4ac))÷2a and x= (-b +(b^2-4ac))÷2aSo x=-2